OpenStudy (anonymous):

solve sec^2x+5tanx=-2

myininaya (myininaya):

sin^2(x)+cos^2(x)=1 so tan^2(x)+1=sec^2(x) so we have tan^2(x)+1+5tanx=-2 tan^2(x)+5tanx+1+2=0 (tanx)^2+5tanx+3=0 let u=tanx so u^2=tan^2(x) or (tanx)^2 u^2+5u+3=0 solve this for u first

OpenStudy (anonymous):

1+tan^2x+5tanx=-2 tan^2x+5tanx+3=0 tanx=(-5+-sqrt(25-12))/2

OpenStudy (anonymous):

hello got it??

OpenStudy (anonymous):

is that the final answer?

myininaya (myininaya):

you want to solve for x take tan inverse of both sides

OpenStudy (anonymous):

so (tan^-1x)^2 +5tan^-1x+3?

myininaya (myininaya):

no

OpenStudy (anonymous):

im confused...

myininaya (myininaya):

tanx=(-5+-sqrt(25-12))/2 take tan inverse (or arctan) of both sides

OpenStudy (anonymous):

u have found the tanx=(-5+-sqrt(25-12))/2 then x=tan^{-1}((-5+-sqrt(25-12))/2)

OpenStudy (anonymous):

thanks for the help myininaya and dipankarstudy

OpenStudy (anonymous):

good luck...